OpenStudy (kenialove):

WILL METAL AND FAN !! Word Problem - A computer store builds custom computers by allowing customers to choose 1 of 4 different CPUs, 1 of 8 hard drives, and 1 of 3 video cards. How many different computers are possible ?

OpenStudy (anonymous):

96. need explanation?

OpenStudy (kenialove):

Yes please

OpenStudy (anonymous):

you just multiply 4 times 8 times 3. So with 1 CPU and changing the hradrives, you can make 8 computers, with 2: 16. just multiply ;)

OpenStudy (kenialove):

I'm sorry can u explain that a little more I didn't quite understand that @arnauvila33

OpenStudy (kenialove):

I understood how u just have to multiply all the substances but I didn't understand the part about with 1 CPU and changing the hard drives

OpenStudy (kenialove):

No wait I think I do understand but wat about the video cards too with that second part

OpenStudy (anonymous):

Sorry for my english, im from Spain so.. alright, look, imagina that you have 2 types of vannilla and 3 of chocolate. and you wanna make a mix ice cream. first you grab one type of vanilla, and you have to add the 3 types of chocolate. That means you get 3 cups full of the same type of vanilla, and on those cups, you put a different type of chocolate. so now u have 3, then you do the same with the other type of vanilla, and have 3 more. 3+3=6 / 2 x 3=6

OpenStudy (anonymous):